# GeoBench: Benchmarking and Analyzing Monocular Geometry Estimation Models

### v1

@@ 5 Benchmarking Geometric Correspondence
(1) Depth-Anything, pretrained with DINOv2 [43] and fine-tuned on K training samples, is comparable to the original DINOv2; Metric3Dv2 pretrained with DINOv2-reg [12] outperforms original DINOv2-reg on ScanNet and inferior to original DINOv2-reg on NAVI dataset. While generative-based models, i.e., Marigold, DepthFM, GenPercept, and Geowizard, are inferior to the original Stable Diffusion [46] model on both datasets.
(2) All models struggle with larger view changes, while generative-based models see a larger drop. In general, monocular geometry estimation models are not 3D-consistent with large viewpoints and thus not yet good enough to encode the 3D structure of the real-world scenario.
@@ 6 Conclusion and Discussion
In this work, we present the first large-scale benchmarking of deterministic and generative geometry estimation foundation models with diverse evaluation datasets. We identify that with a strong pre-train model, either Stable Diffusion or DINOv2, the fine-tuning data quality is a more important factor than fine-tuning data scale and model architecture to achieve generalizable geometry estimation. We believe this benchmarking study can provide strong baselines for unbiased comparisons in geometry estimation studies. Limitations and future works are discussed in the Supp. Mat.

### v2

@@ 5 Benchmarking Geometric Correspondence
1.
Depth-Anything, pretrained with DINOv2 [43] and fine-tuned on K training samples, is comparable to the original DINOv2; Metric3Dv2 pretrained with DINOv2-reg [12] outperforms original DINOv2-reg on ScanNet and inferior to original DINOv2-reg on NAVI dataset. While generative-based models, i.e., Marigold, DepthFM, GenPercept, and Geowizard, are inferior to the original Stable Diffusion [46] model on both datasets.
2.
All models struggle with larger view changes, while generative-based models see a larger drop. In general, monocular geometry estimation models are not 3D-consistent with large viewpoints and thus not yet good enough to encode the 3D structure of the real-world scenario.
@@ 6 Conclusion and Discussion
In this work, we present the first large-scale benchmarking of discriminative and generative geometry estimation foundation models with diverse evaluation datasets. We identify that with a strong pre-train model, either Stable Diffusion or DINOv2, the fine-tuning data quality is a more important factor than fine-tuning data scale and model architecture to achieve generalizable geometry estimation. We believe that this benchmarking study can provide strong baselines for unbiased comparisons in geometry estimation studies. Limitations and future works are discussed in the Supp. Mat.
@@ Appendix A Appendix
Benchmarking: We introduce new benchmarks with diverse scenes and high-quality annotations to overcome limitations in scene diversity and label quality of existing benchmarks.
Findings: The study reveals that discriminative models pre-trained on large data can outperform generative models with less but high-quality synthetic data, suggesting the importance of data quality over quantity and architecture.